# A Survey on Diffusion Models for Recommender Systems

### v1

@@ 3.3.4. Discussion
Generally speaking, adapting diffusion models for content presentation highlights an important evolution in recommender systems. It is no longer limited to selecting and ranking candidate items, but further delves into the user interaction phase and thereby dynamically customizes the content displays of the same items for different users. As shown in Figure 6, we could observe a clear development trend for adapting diffusion models for content presentation from the perspective of generated contents. Earlier studies begin with generating images primarily related to fashion clothing. Then, researchers expand the generative scope and encompass broader range of ad creatives for online advertising. In this circumstance, diffusion models have to generate or edit the image of a certain product under various pre-defined guidelines and intricate instructions, e.g., maintaining the details of the original product, and preserving the coherent and consistent background rendering with the commands from the users or designers. Next, with the emergence of multi-modal foundation models, it evolves beyond the visual modality, and starts to incorporate more generalized and diverse multi-modal data generation, including but not limited to video and audio. In this way, although it imposes much higher demands on the capabilities of generative models, the content generation process itself has become relatively less constrained, allowing for greater creation freedom. This, in turn, is more conducive to dynamically synthesizing contents that aligns with users’ personalized preferences. Finally, we suggest that this progression points toward the ultimate goal of AIGC-driven recommender systems when the correlation and cooperation between the diffusion-based content generation is further deepened with the overall recommendation pipeline. This signals not just the technological progress, but also a profound shift in how humans interact with AI and digital creativity, leading to a new paradigm in the digital content ecosystem. In such a rapidly emerging field of research, there are naturally still many challenges and issues that urgently need to be addressed, which will be discussed in Section 4.
4. Challenges and Future Directions
In this section, we highlight the key challenges in adapting diffusion models to recommender systems. For each challenge, we will further discuss the preliminary efforts done by existing works, as well as other possible solutions.
@@ 4.1. Efficiency and Scalability
  • When adapting diffusion models for data engineering & encoding, the inefficiency issue mainly affects the training phase of recommender systems, since the output of diffusion models (either augmented data or enhanced representations) can be pre-computed and pre-cached to avoid the influence towards the online inference stage. However, the slow generation can limit the training efficiency of downstream recommender models, which typically requires both the large volumes of training data (million- or even billion-level) and the update frequency (from day-level to hour-level or even minute-level). Existing works (Wang et al., 2024b; Liu et al., 2023b) generally adopt the asynchronous update strategy to separate the training of diffusion models and downstream recommenders. In this way, we can reduce the training data volume and relax the update & generation frequency for diffusion models, while maintaining full training data and high update frequency for downstream recommenders.
  • When adapting diffusion models as recommender models, we have to take into account the inference efficiency problem, since the diffusion-based recommenders are directly deployed for real-time online services, where the user request should be responded to within around tens of milliseconds. In this context, preliminary works have been conducted to reduce the inference latency with few-step (Wang et al., 2023b) or even one-step sampling process (Yu et al., 2023a), and possibly combined with an early-stopping strategy (Lin et al., 2024a).
  • When adapting diffusion models for content presentation, the storage overhead should be further considered when tackling the inefficiency issue of diffusion models. As discussed in Section 3.3.4, given the fact that we are currently unable to support real-time online content rendering for each user request, we have to pre-compute and pre-cached the personalized content beforehand. While generating personalized item contents for each single user is impractical due to the storage constraint, existing works (Yang et al., 2024d; Wang et al., 2023a) intend to perform user clustering and pre-cache the group-wise personalized item contents.
@@ 4.2. Integration with Large Language Models
To be specific, LLMs are skilled at interpreting user profiles, queries, and historical interactions by analyzing natural languages, detecting behavior patterns, and discerning subtle semantic cues (Lin et al., 2024b). This integration is particularly important in scenarios where the user’s preferences are complex and dynamic, requiring both a deep understanding of user intent and the ability to generate a wide range of diverse content. In this way, LLMs serve as a bridge between user inputs and the generative capabilities of diffusion models. For instance, by understanding the nuances of a user’s intent, LLMs can convert natural language prompts into meaningful guidance for diffusion-based recommender models, ensuring that the generated recommendations (e.g., ranked item list) or personalized contents (e.g., thumbnails) are not only diverse but also contextually tailored for the user preference. Apart from continuous-space guidance generation, LLMs are also capable of constructing and reformulating individualized prompts for controllable text-to-image generation (Ma et al., 2024e) to improve the quality of content personalization for recommendation. Moreover, the combination of LLMs and diffusion models points towards a more sophisticated and automated reasoning system, which is able to provide more interactive, personalized, and coherent multi-modal recommendations.
@@ 4.4. Digital Copyright and Privacy Preserving
When adapting diffusion models for content presentation, the multi-modal material generation can cause significant digital copyright challenges, particularly related to data ownership, derivative works, and fair use. The diffusion models are often trained on vast datasets that include copyrighted materials, raising the risk of infringement if used without proper permissions. The resulting personalized content may also be considered a derivative work, complicating the issue of who holds ownership—whether it’s the original content creator, the model developer, or the end user. Furthermore, the blending of various copyrighted data sources in multi-modal outputs makes it difficult to trace and attribute the original creators, potentially violating copyright laws. The concept of “fair use” is frequently cited as a defense, but the boundaries between transformative AI-generated content and infringement remain ambiguous for diffusion-based recommender systems. Besides, the adaptation of diffusion models for either data engineering & encoding or as recommender models can also meet the challenge of privacy preserving, where the users’ sensitive personal data should be well preserved from leaking.
To address these challenges, one obvious solution is to source the training data from copyright-free or permissively licensed content, such as public domain works or those under Creative Commons licenses, which can help avoid infringing on the intellectual property rights of creators. Moreover, some cutting-edge research works start to investigate the neural automated watermarking (Liang and Wu, 2023) or text editing strategies (Somepalli et al., 2023) to ensure that the generated content properly attributes the original creators, therefore helping address concerns over authorship. For instance, Liang et al. (2023) propose to employ adversarial samples to add imperceptible perturbations to human-created artworks, which can disturb the training of diffusion models. In this way, the authors establish a powerful toolkit for human creators to protect their artworks from being used without authorization by diffusion-based AIGC applications. In the future, we suggest that the hybrid techniques of neural watermarking (Min et al., 2024), adversarial samples (Costa et al., 2024), federated learning (Li et al., 2021) and differential privacy (Ji et al., 2014) should be vital to mitigate the challenges of digital copyright and privacy preserving in the research field of diffusion-based recommender systems.
@@ 5. Conclusion
In this survey, we provide a comprehensive review of the research efforts on adapting diffusion models to recommender systems. We systematically classify existing research works into three primary categories: (1) diffusion for data engineering & encoding, which focuses on data augmentation and representation enhancement; (2) diffusion as recommender models, which employs diffusion models to directly estimate user preferences and rank items; and (3) diffusion for content presentation, which leverages diffusion models to generate personalized content such as fashion and advertisement creatives. We also give detailed discussion about the core characteristics of the adapting diffusion models for recommendation, and further identify key challenges and future directions for exploration. We hope this survey could serve as a foundational roadmap for researchers and practitioners to advance recommender systems through the innovative application of diffusion models.

### v2

@@ 3.3.4. Discussion
Generally speaking, adapting diffusion models for content presentation highlights an important evolution in recommender systems. It is no longer limited to selecting and ranking candidate items, but further delves into the user interaction phase and thereby dynamically customizes the content displays of the same items for different users. As shown in Figure 6, we could observe a clear development trend for adapting diffusion models for content presentation from the perspective of generated contents. Earlier studies begin with generating images primarily related to fashion clothing. Then, researchers expand the generative scope and encompass a broader range of ad creatives for online advertising. In this circumstance, diffusion models have to generate or edit the image of a certain product under various pre-defined guidelines and intricate instructions, e.g., maintaining the details of the original product, and preserving the coherent and consistent background rendering with the commands from the users or designers. Next, with the emergence of multi-modal foundation models, it evolves beyond the visual modality, and starts to incorporate more generalized and diverse multi-modal data generation, including but not limited to video and audio. In this way, although it imposes much higher demands on the capabilities of generative models, the content generation process itself has become relatively less constrained, allowing for greater creation freedom. This, in turn, is more conducive to dynamically synthesizing content that aligns with users’ personalized preferences. Finally, we suggest that this progression points toward the ultimate goal of AIGC-driven recommender systems when the correlation and cooperation between the diffusion-based content generation is further deepened with the overall recommendation pipeline. This signals not just the technological progress, but also a profound shift in how humans interact with AI and digital creativity, leading to a new paradigm in the digital content ecosystem. In such a rapidly emerging field of research, there are naturally still many challenges and issues that urgently need to be addressed, which will be discussed in Section 4.
4. Future Directions and Challenges
In this section, we cast detailed discussions on future directions on the research of adapting diffusion models to recommender systems. We would highlight the key challenges, and further discuss the preliminary efforts done by existing works, as well as other possible solutions.
@@ 4.1. Efficiency and Scalability
  • When adapting diffusion models for data engineering & encoding, the inefficiency issue mainly affects the training phase of recommender systems, since the output of diffusion models (either augmented data or enhanced representations) can be pre-computed and pre-cached to avoid the influence towards the online inference stage. However, the slow generation can limit the training efficiency of downstream recommendation models, which typically requires both the large volumes of training data (million- or even billion-level) and the update frequency (from day-level to hour-level or even minute-level). Existing works (Wang et al., 2024b; Liu et al., 2023b) generally adopt the asynchronous update strategy to separate the training of diffusion models and downstream recommenders. In this way, we can reduce the training data volume and relax the update & generation frequency for diffusion models, while maintaining full training data and high update frequency for downstream recommenders.
  • When adapting diffusion models as recommendation models, we have to take into account the inference efficiency problem, since the diffusion-based recommenders are directly deployed for real-time online services, where the user request should be responded to within around tens of milliseconds. In this context, preliminary works have been conducted to reduce the inference latency with few-step (Wang et al., 2023b) or even one-step sampling process (Yu et al., 2023a), and possibly combined with an early-stopping strategy (Lin et al., 2024a).
  • When adapting diffusion models for content presentation, the storage overhead should be further considered. As discussed in Section 3.3.4, given the fact that we are currently unable to support high-resolution real-time online content rendering for each user request (i.e., inference inefficiency), we have to pre-compute and pre-cached the personalized content beforehand. The operations of pre-storing individualized contents could then meet the storage overhead, i.e., we have to expand the space complexity from $O(M)$ for $M$ items to $O(MN)$ for $M\times N$ user-item pairs. While generating personalized item contents for each single user is impractical due to such a storage constraint, existing works (Yang et al., 2024d; Wang et al., 2023a) intend to perform user clustering and pre-cache the group-wise personalized item contents.
@@ 4.2. Integration with Large Language Models
To be specific, LLMs are skilled at interpreting user profiles, queries, and historical interactions by analyzing natural languages, detecting behavior patterns, and discerning subtle semantic cues (Lin et al., 2024b). This integration is particularly important in scenarios where the user’s preferences are complex and dynamic, requiring both a deep understanding of user intent and the ability to generate a wide range of diverse content. In this way, LLMs serve as a bridge between user inputs and the generative capabilities of diffusion models. For instance, by understanding the nuances of a user’s intent, LLMs can convert natural language prompts into meaningful guidance for diffusion-based recommendation models, ensuring that the generated recommendations (e.g., ranked item list) or personalized contents (e.g., thumbnails) are not only diverse but also contextually tailored for the user preference. Apart from continuous-space guidance generation, LLMs are also capable of constructing and reformulating individualized prompts for controllable text-to-image generation (Ma et al., 2024e) to improve the quality of content personalization for recommendation. Moreover, the combination of LLMs and diffusion models points towards a more sophisticated and automated reasoning system, which is able to provide more interactive, personalized, and coherent multi-modal recommendations.
@@ 4.4. Digital Copyright and Privacy Preserving
When adapting diffusion models for content presentation, the multi-modal material generation can cause significant digital copyright challenges, particularly related to data ownership, derivative works, and fair use. The diffusion models are often trained on vast datasets that include copyrighted materials, raising the risk of infringement if used without proper permissions. The resulting personalized content may also be considered a derivative work, complicating the issue of who holds ownership—whether it’s the original content creator, the model developer, or the end user. Furthermore, the blending of various copyrighted data sources in multi-modal outputs makes it difficult to trace and attribute the original creators, potentially violating copyright laws. The concept of “fair use” is frequently cited as a defense, but the boundaries between transformative AI-generated content and infringement remain ambiguous for diffusion-based recommender systems. Besides, the adaptation of diffusion models for either data engineering & encoding or as recommendation models can also meet the challenge of privacy preserving, where the users’ sensitive personal data should be well preserved from leaking.
To address these challenges, one obvious solution is to source the training data from copyright-free or permissively licensed content, such as public domain works or those under Creative Commons licenses, which can help avoid infringing on the intellectual property rights of creators. Moreover, some cutting-edge research works start to investigate the neural automated watermarking (Liang and Wu, 2023) or text editing strategies (Somepalli et al., 2023) to ensure that the generated content properly attributes the original creators, therefore helping address concerns over authorship. For instance, Liang et al. (2023) proposes to employ adversarial samples to add imperceptible perturbations to human-created artworks, which can disturb the training of diffusion models. In this way, the authors establish a powerful toolkit for human creators to protect their artworks from being used without authorization by diffusion-based AIGC applications. In the future, we suggest that the hybrid techniques of neural watermarking (Min et al., 2024), adversarial samples (Costa et al., 2024), federated learning (Li et al., 2021) and differential privacy (Ji et al., 2014) should be vital to mitigate the challenges of digital copyright and privacy preserving in the research field of diffusion-based recommender systems.
@@ 5. Conclusion
In this survey, we provide a comprehensive review of the research efforts on adapting diffusion models to recommender systems. We systematically classify existing research works into three primary categories: (1) diffusion for data engineering & encoding, which focuses on data augmentation and representation enhancement; (2) diffusion as recommendation models, which employs diffusion models to directly estimate user preferences and rank items; and (3) diffusion for content presentation, which leverages diffusion models to generate personalized content such as fashion and advertisement creatives. We also give detailed discussion about the core characteristics of the adapting diffusion models for recommendation, and further identify key challenges and future directions for exploration. We hope this survey could serve as a foundational roadmap for researchers and practitioners to advance recommender systems through the innovative application of diffusion models.